# RadGPT: Constructing 3D Image-Text Tumor Datasets

### v1

@@ 3.1.2 Segment Tumors and Measure Like Radiologists
Standardized tumor measurement is key for cancer prognosis and treatment planning [47, 38]. Human-made radiology reports commonly use the World Health Organization (WHO) tumor measurement standard, which measures tumors with two diameters: the longest tumor diameter in any axial plane ($D$), and its perpendicular diameter in the same plane ($d$) [47]. Thus, to measure tumors like radiologists in our structured reports, RadGPT first needs tumor segmentations. AbdomenAtlas 3.0 presents ground-truth segmentations of liver, kidney and pancreas tumors. For fully-automated report generation, RadGPT segments tumors with DiffTumor, a public AI model [14, 15]. From segmentations, RadGPT extracts tumor measurements using a deterministic algorithm, which we designed to implement the WHO measuring standard (Alg. 2). Additionally, our structured reports present tumor and organ volume and attenuation (HU values), both extracted from segmentation. AbdomenAtlas 3.0 has ground-truth segmentations for multiple organs [^3] AbdomenAtlas 3.0 is the first dataset with per-voxel annotations for the blood vessels key for pancreatic tumor staging: the celiac axis (CA), superior mesenteric artery (SMA), superior mesenteric vein (SMV), common hepatic artery (CHA), and portal vein. It also has per-voxel annotations for other 22 structures important for cancer detection/staging: liver tumors, kidney tumors, pancreas tumors, liver, kidney, pancreas, spleen, adrenal glands, stomach, duodenum, bile duct, intestines, aorta, and postcava. . For fully-automated report generation, RadGPT segments organs with an nnU-Net trained in AbdomenAtlas 1.1 [7, 37]. Organ volumes allows our structured reports to diagnose enlarged organs, and attenuation allows diagnosing fatty liver (average HU $<40$ [32]) and pancreas (pancreas-to-spleen attenuation $<0.7$ [21])—a condition related to diabetes and pancreatic cancer [21]. Meanwhile, tumor attenuation helps diagnose tumor type.
@@ 3.1.3 Stage Pancreatic Cancer using Segmentation
Cancer T stage indicates the tumor size and its interaction with nearby structures. It is vital for surgical planning and survival, especially for pancreatic adenocarcinoma (PDAC), a highly aggressive cancer [1]. However, staging is time-consuming for radiologists. To provide PDAC T stage in our structured reports, we measure the tumor (§3.1.2) and analyze its interaction with nearby blood vessels (SMA, CHA, CA, SA3) [1]. Thus, RadGPT first segments these vessels and tumors, using an nnU-Net trained on private data and DiffTumor, respectively. Then, a deterministic algorithm uses the segmentations to measure the tumor-vessel angle of contact (Alg. 3). Large angles ($>$180) make surgical removal challenging, increasing stage. For interpretability, the reports justify stages with tumor size and tumor-vessel degree of contact. AbdomenAtlas 3.0 is first public CT dataset with PDAC T stage annotations.
Large Tumors $>$2 cm
Liver Tumor (HCC) Kidney Tumor (RCC) Pancreas Tumor (PDAC) CRC Liver Metastases
Model Sensitivity (%) Specificity (%) Sensitivity (%) Specificity (%) Sensitivity (%) Specificity (%) Sensitivity (%) Specificity (%)
CT2Rep 0.0 (0/301) 100.0 (244/244) 10.0 (22/219) 98.0 (239/244) 3.8 (4/105) 96.7 (236/244) 9.4 (2/21) 98.1 (155/158)
M3D 14.6 (44/301) 88.9 (217/244) 9.1 (20/219) 92.6 (226/244) 4.8 (5/105) 96.3 (235/244) 12.1 (7/58) 85.4 (135/158)
RadGPT (ours) 89.4 (269/301) 73.4 (179/244) 97.3 (213/219) 78.3 (191/244) 91.4 (96/105) 76.6 (187/244) 100.0 (58/58) 70.3 (111/158)
Small Tumors $\leq$ 2 cm
CT2Rep 0.0 (0/142) 100.0 (244/244) 4.0 (2/50) 98.0 (239/244) 2.1 (8/385) 96.7 (236/244) 3.4 (2/58) 98.1 (155/158)
M3D 6.3 (9/142) 88.9 (217/244) 2.0 (1/50) 92.6 (226/244) 2.8 (11/385) 96.3 (235/244) 14.3 (3/21) 85.4 (135/158)
RadGPT (ours) 79.6 (113/142) 73.4 (179/244) 92.0 (46/50) 78.3 (191/244) 76.9 (296/385) 76.6 (187/244) 100.0 (21/21) 70.3 (111/158)
Table 2: In tumor detection, fully-automated reports by RadGPT surpass reports created by end-to-end report generation models. Here, we use RadGPT as a fully-automated two-stage pipeline (Figure 2). RadGPT surpasses the two available end-to-end 3D CT report generation models: M3D [4], the state-of-the-art in abdominal CT report generation, and CT2Rep [23], the state-of-the-art in chest CT report generation, which we adapted to abdominal CT scans. We performed out-of-distribution (OOD) evaluation at UCSF, a hospital not seen in training. The results indicate that per-voxel segmentation (step 1 in the RadGPT pipeline) may strongly benefit report generation.
@@ 3.2 Creating Narrative Reports
Structured reports use rigid templates to improve clarity and clinical decision-making [1]. However, rigid templates may conflict with the reporting style of an institution. Thus, RadGPT can create narrative reports that mimic the style of a target institution. In AbdomenAtlas 3.0, they mimic human-made reports at UCSF (Figure 1). The narrative reports are created through style adaptation with in-context learning: we provide a pre-trained LLM [^4] Llama-3.1 with 70B parameters and AWQ quantization [20] with a structured report and 10 human-made reports from the target institution, and the LLM adapts the structured report to the style of the human-made reports. We ask the LLM not to change any diagnosis. Thus, narrative reports contain all the detailed information found in structured reports (§3.1).
However, the style of human-made reports varies significantly with its diagnoses. E.g., pancreatic tumor reports differ from liver tumor or pneumonia reports [48]. Thus, we verify diagnoses to give the LLM human-made reports with the correct style. We first use an LLM to categorize human-made reports according to their tumors (liver, pancreas, kidney, or none). Then, when adapting a structured report into narrative, we give the LLM human-made reports with the same tumor as the structured report (e.g., liver).
After adapting a structured report to a narrative report, the LLM performed a quality check. It extracted diagnoses and quantitative information (e.g., tumor size and stage) from both reports and checked for consistency. We prompted the LLM to correct in the narrative report any information diverging from the structured report, and to remove any diagnosis not present in the structured report.
3.3 Creating Human-AI Fusion Reports
The structured and narrative reports in AbdomenAtlas 3.0 focus on tumors. Our reports can quantify multiple tumors in a CT scan so they are much more detailed than human-made reports (§4.5). However, human-made reports cover multiple diagnoses unrelated to tumors. To get the best of both worlds, RadGPT prompts a zero-shot LLM4 to fuse the details in structured reports with the many diagnoses in human-made reports or clinical notes (Figures 11 and 1), generating fusion reports. AbdomenAtlas 3.0 has 240 of them: 209 used clinical notes for TotalSegmentator CT scans [4], and 31 used notes from collaborating radiologists.
@@ 3.4 Evaluating Diagnoses in AI-Made Reports
We propose an automatic strategy to evaluate the clinical utility of AI-made reports. Specifically, it evaluates whether the diagnoses in the reports are correct. We prompt an LLM4 (see prompts in Appendix B.3) to identify whether reports mention liver, kidney, or pancreatic tumors. Then, we convert the LLM answers into categorical labels. We create labels for both AI-made reports and human-made reports (ground-truth). By comparing them, we calculate tumor detection sensitivity and specificity for the AI-made reports. Besides producing clinically relevant metrics (§4.4), this evaluation strategy is practical: using zero-shot inference, it does not need fine-tuning and is easily adaptable to multiple hospitals. Also, with a simple prompt modification, it can evaluate multiple diagnoses beyond tumors.
@@ 4 Experiment & Result
In AbdomenAtlas 3.0, RadGPT leverages ground-truth tumor annotations, ensuring reports match the accuracy of these annotations. However, this section evaluates RadGPT as a fully automated two-stage pipeline, without manual annotations, as shown in Figure 2. Stage I: AI models segment tumors and organs, using DiffTumor [14, 15] for tumors and SuPreM [37] for organs. Stage II: RadGPT converts these segmentations into structured reports. We compare our two-stage approach (RadGPT) to end-to-end report generation approaches (M3D [4] and CT2Rep [23]). To ensure realistic evaluation [7, 6], we test on out-of-distribution hospitals not included in the training data.
4.1 RadGPT Enhances Tumor Detection
To enable automated evaluation on a large test dataset, we use an LLM (Llama-3.1) to assess the reports generated by RadGPT (§3.4). First, we verify the LLM’s ability to determine whether a report indicates the presence of a tumor. As shown in Figure 3, the zero-shot LLM achieves 96% accuracy on this task, demonstrating its reliability for evaluating tumor detection sensitivity and specificity.
The LLM evaluation showed that the reports generated by RadGPT strongly surpassed existing abdominal CT report generation models. Table 2 compares RadGPT reports to M3D [4], the only public AI model for abdominal CT report generation, and to CT2Rep [23], the first AI model for report generation in chest CT scans. Here, we adapted CT2Rep to the abdomen region (Appendix B.1). Both CT2Rep and M3D had difficulty detecting tumors (low sensitivity), and RadGPT strongly outperformed them for small and large tumors in the liver, pancreas, and kidneys. This performance difference shows the benefits of using segmentation to improve report generation: DiffTumor produces accurate tumor segmentations, which RadGPT translates into reports. By releasing AbdomenAtlas 3.0, the first abdominal CT dataset with triplets of CT scans, reports, and per-voxel annotations, our objective is to catalyze further research on segmentation-driven report generation.
RadGPT is the first AI model to perform cancer staging on abdominal CT. Figure 3 shows the result of RadGPT on the staging of pancreatic adenocarcinoma. The RadGPT fully-automated reports achieved accuracy of 71.43% in determining tumor T stages 1 to 3. The results show that AI is a promising tool for assisting cancer staging, a key but time-consuming task for radiologists.
@@ 4.1 RadGPT Enhances Tumor Detection
Figure 3: Confusion matrices. (a) A zero-shot LLM (Llama 3.1) has 96% accuracy and 0.953 F1-score in determining whether radiology reports indicate tumors. Thus, the LLM can accurately calculate tumor detection sensitivity and specificity for AI-made reports (§3.4). The LLM accuracy rivals established labelers, like those in CheXpert [28] and CheX-ray14 [61]. Results were manually evaluated by radiologists. They used 447 randomly selected UCSF reports including kidney, pancreas, and liver tumors. (b) PDAC staging confusion matrix for RadGPT, the first public DNN for staging abdominal CT tumors. Results consider a private dataset with ground-truth tumor stage annotations ($N$=42).
@@ 4.2 RadGPT Accurately Measures Tumor Size
An expert radiologist manually evaluated structured reports generated by RadGPT. He analyzed each reported tumor, evaluating its measurement and checking if the tumor is a false-positive (tumor not present in the CT volume) or a true-positive (present). The radiologist deemed 75.6% of the tumors reported by RadGPT true-positives, and 93.8% of them were correctly measured (Table 3). RadGPT only made measuring mistakes for pancreatic tumors (PDAC), but even the radiologist could not measure 3 PDACs.
Tumor # Correct Tumor Detection # Tumor Detection # Correct Tumor Size Tumor Detection Precision (%) Size Accuracy (%)
HCC 12 13 13 92.3 100.0
PDAC 8 16 6 50.0 75.0
RCC 11 12 11 91.7 100.0
Total 31 41 30 75.6 93.8
Table 3: RadGPT achieved 75.6% tumor detection precision and 93.8% tumor size measurement accuracy, according to manual evaluation. A radiologist evaluated the reports RadGPT created for 23 CT scans from UCSF. A reported tumor measurement was considered correct if it deviated by 10% or less from the radiologist’s measurement (both use the WHO measuring standard [47]). As manual evaluation is time-consuming, the radiologist evaluated 23 reports. Using an LLM for automatically evaluating tumor measurements is challenging: it requires pairing tumors in AI-made reports and ground-truth reports. E.g., if both reports mention 3 tumors in liver sub-segment 2, it is hard to understand which measurements we should compare.

### v2

@@ 3.1.2 Measure Tumors Like Radiologists
Radiologists commonly measure tumors using use the World Health Organization (WHO) standard, which provides two diameters: the longest tumor diameter in any axial plane ($D$), and its perpendicular diameter in the same plane ($d$) [40]. Standardization of measurements is key for accurate cancer prognosis and treatment planning [40, 32]. Thus, RadGPT also uses the WHO standard, measuring tumors like radiologists. AbdomenAtlas 3.0 presents radiologist-revised segmentations of liver, kidney and pancreas tumors. From segmentations, RadGPT extracts tumor measurements using a deterministic algorithm that implements the WHO standard (Sup. Alg. 2). Besides diameters, our structured reports present tumor & organ volume and attenuation (HU values), also extracted from segmentation [^1] AbdomenAtlas 3.0 is the first dataset with per-voxel annotations for the blood vessels key for pancreatic tumor staging: the celiac axis (CA), superior mesenteric artery (SMA), superior mesenteric vein (SMV), common hepatic artery (CHA), and portal vein. These annotations were produced by an nnU-Net trained in private data, and revised by radiologists (Appendix C). AbdomenAtlas 3.0 also has per-voxel annotations for other 22 structures important for cancer detection/staging: liver tumors, kidney tumors, pancreas tumors, liver, kidney, pancreas, spleen, adrenal glands, stomach, duodenum, bile duct, intestines, aorta, and postcava. . Using volumes, our reports diagnose enlarged organs, and attenuation diagnoses fatty liver (average HU <40<40 [28]) and pancreas (pancreas-to-spleen attenuation <0.7<0.7 [18])—a condition related to diabetes and pancreatic cancer [18]. Meanwhile, tumor attenuation helps identify tumor type.
Figure 3: Automated T staging. Our RadGPT first segments the tumor and key vascular structures from CT scans, then measures tumor size and blood vessel contact angle to automatically assign T stage and resectability. If the tumor-vessel contact angle surpasses 180 degrees, the tumor becomes unresectable (T stage 4).
@@ 3.1.3 Stage Pancreatic Cancer using Segmentation
Tumor T‑stage summarizes tumor size and relationship to nearby structures. It is key for surgical planning and survival, especially for pancreatic adenocarcinoma (PDAC), an aggressive cancer [1]. However, staging is time-consuming. As shown in Figure 3, for PDAC staging, radiologists must measure tumors (§3.1.2) and analyze its interaction with blood vessels (SMA, CHA, CA, SA) [1]. Accordingly, RadGPT first segments vessels and tumors (using nnU-Net and DiffTumor1) and radiologists revise segmentations (Appendix C). Then, a deterministic algorithm uses the revised segmentations to measure the tumor-vessel contact angle (Sup. Alg. 3). Large angles (>>180) make surgery difficult, increasing stage. For interpretability, reports justify stages with tumor size and tumor-vessel degree of contact, and our deterministic algorithm faithfully implements the guidelines radiologists use to stage PDAC [1]. AbdomenAtlas 3.0 is first public dataset with PDAC T stage labels.
Internal validation on the test set of AbdomenAtlas 3.0 (IID)
pancreatic tumor (%) kidney tumor (%) liver tumor (%)
Model Sen. ($\leq$2 cm) Sen. (>>2 cm) Spec. Sen. ($\leq$2 cm) Sen. (>>2 cm) Spec. Sen. ($\leq$2 cm) Sen. (>>2 cm) Spec.
CT-CHAT [20] 66.7 51.9 61.2 31.1 32.8 74.2 5.7 3.2 94.7
CT2Rep [21] 0.0 0.0 92.5 36.5 39.3 70.4 35.8 49.2 70.4
M3D [4] 0.0 7.4 97.2 8.1 16.4 84.1 9.4 12.7 86.0
Merlin [8] 33.3 51.9 71.8 28.4 45.9 86.6 30.2 41.3 95.9
RadFM [54] 0.0 0.0 99.9 3.7 6.3 95.6 3.3 5.7 93.9
RadGPT (ours) 66.7 81.5 93.2 54.8 93.3 51.8 39.6 96.8 64.4
External validation on unseen hospital—UCSF (OOD)
Model Sen. ($\leq$2 cm) Sen. (>>2 cm) Spec. Sen. ($\leq$2 cm) Sen. (>>2 cm) Spec. Sen. ($\leq$2 cm) Sen. (>>2 cm) Spec.
CT-CHAT [20] 27.5 N/A 73.1 24.3 29.7 74.6 5.2 4.2 94.0
CT2Rep [21] 2.1 N/A 96.7 4.0 10.0 98.0 0.0 0.0 100.0
M3D [4] 3.3 N/A 97.9 14.8 13.1 86.3 10.7 17.3 87.3
Merlin [8] 7.5 N/A 100.0 8.1 9.2 100.0 9.1 19.2 100.0
RadFM [54] 0.0 N/A 100.0 7.5 6.8 90.9 10.9 11.1 85.0
RadGPT (ours) 76.9 N/A 76.6 92.0 97.3 78.3 79.6 89.4 73.4
Table 2: In tumor detection, fully-automated reports by RadGPT surpass reports created by end-to-end report generation models. We use RadGPT as a fully-automated segmentation-assisted report generation model (Figure 2). The results indicate that per-voxel segmentation (step 1 in the RadGPT pipeline) may strongly improve report generation. We tested out-of-distribution (OOD) at UCSF, a hospital not seen in training, and in-distribution (IID). In the IID set, the ground-truth contained 9 small and 27 large pancreatic tumors, 74 small and 61 large kidney tumors, and 53 small and 63 large liver tumors, with 890, 791, and 810 negatives, respectively. In OOD, we have 385 (small) and 0 (large) for pancreas, 50 (small) and 219 (large) for kidney, and 142 (small) and 301 (large) for liver, with 244 negatives for each organ. Decision thresholds are analyzed in 15. While other methods were evaluated zero-shot, CT-CHAT, CT2Rep and Merlin were trained in AbdomenAtlas 3.0, giving them an advantage in the IID dataset. To compute sensitivity and specificity, we used our proposed diagnostic evaluation (§3.4): an LLM extracted binary tumor presence labels per-organ, and we compared the labels for AI-made reports and ground-truth human-made reports. LLM label extraction accuracy is 96% (Figure 4). Table 5 provides additional metrics (BLEU, ROUGE, BERT, RadGraph-F1), showing they are usually sensible to variations in report style, unlike our diagnostic evaluation.
@@ 3.2 Creating Narrative Reports
Structured reports use rigid templates to improve clarity and clinical decision-making [1]. However, rigid templates may conflict with the reporting style of an institution. Thus, RadGPT can create narrative reports that mimic the style of a target institution. In AbdomenAtlas 3.0, they mimic human-made reports at a major US hospital (Figure 1). The narrative reports are created through style adaptation with in-context learning: we provide a pre-trained LLM (Llama-3.1 70B, AWQ quantization [17]) with a structured report and 10 human-made reports from the target institution, and the LLM adapts the structured report to the style of the human-made reports. We ask the LLM not to change diagnoses or details. Thus, narrative reports contain all the detailed information from structured reports (§3.1).
However, style of human-made reports varies with diagnoses. E.g., pancreatic tumor differ from liver tumor reports [41]. Thus, we verify diagnoses to give the LLM example reports with the correct style. First, another LLM categorizes human-made reports according to tumors (liver, pancreas, kidney, none). Then, when adapting a structured report to narrative, the first LLM receives example human-made reports with the same tumor as the structured report.
After adapting a structured report into a narrative report, the LLM performed a quality check. It extracted diagnoses and quantitative information (e.g., tumor size and stage) from both reports and checked for consistency. We prompted the LLM to correct in the narrative report any information diverging from the structured report, and to remove any diagnosis not present in the structured report.
3.3 Creating Enhanced Human Reports
Like most abdominal CT datasets, AbdomenAtlas 3.0 focuses on tumors—as cancer is a major cause of death. Our reports can precisely measure and analyze multiple tumors in a CT, while human-made reports usually measure the largest tumors only (§4.3). However, human-made reports cover multiple diagnoses unrelated to tumors. To combine their strengths, RadGPT prompts the zero-shot LLM (Llama 3.1 70B AWQ) to fuse the details in structured reports with the many diagnoses in human-made reports/clinical notes (Figures 12 and 1), generating enhanced human reports. AbdomenAtlas 3.0 has 240 of them: 209 used clinical notes for TotalSegmentator CT scans [4], and 31 used notes from our radiologists. They span 66 diagnoses.
@@ 3.4 Evaluating Diagnoses in AI-Made Reports
We propose a new strategy to evaluate the clinical utility of AI-made reports: a straightforward, LLM-based diagnostic evaluation. First, we prompt a zero-shot LLM (Llama 3.1 70B AWQ, prompts in §B.4) to identify in which organ the report mentions tumors. Then, we convert the LLM answer into categorical labels. We compare labels for AI-made and human-made reports (ground-truth) to calculate tumor detection sensitivity and specificity. This evaluation strategy is scalable and practical: with zero-shot inference, it does not need fine-tuning and is easily adaptable to multiple hospitals. Importantly, our strategy produces clinically relevant metrics (detection sensitivity / specificity), which are easy to interpret by clinicians. Here, we limit our evaluation strategy to tumor detection. However, it can be expanded to evaluate other relevant clinical information and diseases beyond tumors—with simple prompt modifications.
@@ 4 Experiment & Result
Figure 4: Confusion matrices. (a) A zero-shot LLM (Llama 3.1 70B AWQ) has 96% accuracy, 0.953 F1-score in determining if radiology reports show tumors. Thus, the LLM can accurately calculate tumor detection sensitivity and specificity for AI-made reports (§3.4). LLM’s accuracy rivals established labelers, as those in CheXpert [24] and CheX-ray14 [50]. Results were manually evaluated by radiologists on 447 reports with kidney, pancreas, and liver tumors. (b) PDAC staging confusion matrix for RadGPT, the first public AI for staging abdominal CT tumors. Results on a private dataset with ground-truth tumor stage annotations (NN=42).
We randomly selected 10% of AbdomenAtlas 3.0 as a test set, where we evaluated 6 CT report generation models: CT2Rep [21], M3D [4], CT-CHAT [20], Merlin [8], RadFM [54] and RadGPT—as baselines for future work. In AbdomenAtlas 3.0, RadGPT transforms radiologist-revised tumor masks into reports. In this section, we evaluate RadGPT as a fully-automated, segmentation-assisted method, without radiologist revision (Figure 2). We have both AI trained on AbdomenAtlas 3.0 (CT2Rep and CT-CHAT, see Appendix B.1 for training details) and those trained on abdominal CTs in other works (M3D, Merlin, DiffTumor inside RadGPT). To ensure realistic evaluation [6, 5], we evaluate on the AbdomenAtlas 3.0 test set and on a private out-of-distribution (OOD) dataset, from the University of California San Francisco hospital (UCSF, California, USA) never seen by any AI in training.
Zero-shot LLMs can accurately evaluate report generation. For automated evaluation on a large test dataset, we will use an LLM (Llama-3.1) to assess the reports generated by 6 AI models (§3.4). Before that, radiologists verified the LLM’s ability to determine whether a report indicates tumors or not. They read the zero-shot LLM answers for 447 different reports, verifying that it achieved 96% accuracy (Figure 4). Results demonstrate the LLM reliability in evaluating tumor detection, per-organ.
Segmentation can assist report generation models. The LLM-based evaluation, (Table 2) showed that the reports generated by RadGPT strongly surpassed the other abdominal CT report generation models, especially in the OOD test set (unseen hospital) [^2] As RadGPT narrative and structured reports match in diagnostic accuracy we present only one result for RadGPT.. End-to-end trained methods had difficulty detecting tumors in the OOD dataset (low sensitivity), and RadGPT strongly outperformed them for small and large tumors in the liver, pancreas, and kidneys. This performance difference shows the benefits of using segmentation to improve report generation: DiffTumor produces accurate tumor segmentations, which RadGPT translates into reports. By releasing AbdomenAtlas 3.0, the first abdominal CT dataset with triplets of CT scans, reports, and per-voxel annotations, our objective is to catalyze further research on segmentation-assisted report generation.
RadGPT is the first public AI model to perform cancer staging on abdominal CT. Figure 4 shows the performance of RadGPT for staging of pancreatic adenocarcinoma. RadGPT fully-automated reports achieved accuracy of 71.43% in determining tumor T stages 1 to 3. The results show that AI is a promising tool for assisting cancer staging, a key but time-consuming task for radiologists. Still, these fully-automatic results show radiologist revision is essential to ensure staging accuracy in AbdomenAtlas 3.0.
@@ 4.1 RadGPT Accurately Measures Tumor Size
liver tumor pancreatic tumor kidney tumor
Detection Precision (%) 92.3 (12/13) 50.0 (8/16) 91.7 (11/12)
Size Accuracy (%) 100.0 (12/12) 75.0 (6/8) 100.0 (11/11)
Table 3: RadGPT has 75.6% tumor detection precision and 93.5% tumor measurement accuracy. A radiologist manually evaluated reports RadGPT created for 23 external test CTs (UCSF). A tumor measurement was considered correct if it deviated by $\leq$10% from the radiologist’s measurement (both use the WHO measuring standard [40]). As evaluation is time-consuming, the radiologist evaluated 23 reports. Using an LLM for automatically evaluating tumor measurements is challenging: it requires pairing tumors in AI-made reports and ground-truth reports.
An expert radiologist manually evaluated structured reports generated by RadGPT. He analyzed each reported tumor, evaluating its measurement and checking if the tumor is a false-positive (tumor not present in the CT volume) or a true-positive (present). The radiologist deemed 75.6% of the tumors reported by RadGPT true-positives, and 93.5% of them were correctly measured (Table 3). RadGPT only made measuring mistakes for pancreatic tumors (PDAC), but even the radiologist could not measure 3 PDACs.